OpenStudy (anonymous):

A surveyor has determined that a mountain is h = 2470 ft high. From the top of the mountain he measures the angles of depression to two landmarks at the base of the mountain and finds them to be 42° and c = 38°. The angle between the lines of sight to the landmarks is 68°. Calculate the distance between the two landmarks.

OpenStudy (anonymous):

This is an image that is with the problem and these words were included in the description: (Observe that these are the same as the angles of elevation from the landmarks as shown in the figure below.)

OpenStudy (anonymous):

look at the third triangle and find two sides of it using the height and the angles provided

OpenStudy (anonymous):

Using these:\[\frac{ 2470 }{ \sin38 }=\frac{ x }{\sin52}\] x=3161.4 \[\frac{ 2470 }{ \sin42 }=\frac{ x }{\sin16}\] x=1017.5 Then I found the angle in between to be 148. Then using law of cosines I did:\[a^2 = 3161.4^2 + 1017.5^2 - 2(3161.4)(1017.5)\cos(148)\] This resulted in a= 4060 (I only needed it as an integer) This answer was incorrect. Where did I go wrong?

OpenStudy (anonymous):

sin 32=.... sin 42=....

OpenStudy (anonymous):

you could then find the two sides of the angle 68 and then use the cos law

OpenStudy (anonymous):

I think I must not be following. I still got 4060.2 =/

OpenStudy (anonymous):